# Words and Action: Modeling Linguistic Leadership in #BlackLivesMatter Communities

### v1

@@ Abstract
In this project we describe a method of modeling semantic leadership across a set of communities associated with the #BlackLivesMatter movement, which has been informed by qualitative research on the structure of social media and Black Twitter in particular. We describe our bespoke approaches to time-binning, community clustering, and connecting communities over time, as well as our adaptation of state-of-the-art approaches to semantic change detection and semantic leadership induction. We find substantial evidence of the leadership role of BLM activists and progressives, as well as Black celebrities. We also find evidence of the sustained engagement of the conservative community with this discourse, suggesting an alternative explanation for how we arrived at the present moment, in which “anti-woke” and “anti-CRT” bills are being enacted nation-wide.
@@ Introduction
Battles over social and political change are always waged with words and action–one need only look to the “Stop W.O.K.E. Act,” recently signed into Florida law, which prohibits the discussion of race and racism in public school and university classrooms, for evidence of how words index larger concepts and debates (?). While this particular bill was designed to censor discourse, ample evidence exists of how words can open up conversations, introduce new ideas into public consciousness, and impact social behavior and government policy (??). In this paper, we turn to one of the most visible contemporary examples of the intersection of words and action–the Black Lives Matter movement–to explore how words enter into a specific sociopolitical conversation, how they transform as that conversation evolves, and who is responsible for carrying forward those new or changed meanings.
@@
Figure 1: Semantic leadership in #BlackLivesMatter Nodes represent the twelve main communities of the BLM network. The thickness of the edges are proportional to the number of word changes shared by each pair.
@@
While there exists abundant research on Black Lives Matter (BLM), both as a hashtag (?????) and a movement (???), there are still open questions about how precisely BLM was able to “push the mainstream public sphere on issues of social progress” (?, p. xxvii). This is largely due to the unique community structure of Twitter, in which communities are defined not by individual users but by hashtags, and by the conversations among users that hashtags enable (????). Operationalizing these medium-specific theories about community formation on Twitter, as well as about “Twitter time,” we present a multipart model that allows us to 1) detect the latent communities associated with the #BlackLivesMatter hashtag and link them over time; and 2) detect semantic changes associated with each community using word embeddings. Together with a measure of semantic leadership (?), we are able to induce a semantic leadership network (see Figure 1), allowing us to identify the communities responsible for introducing new or changed word meanings into the BLM network and the communities responsible for adopting (or co-opting) them.
@@
We find that the community of BLM activists plays an outsized role in introducing semantic changes into the network, and that these changes are most consistently taken up by the conservative community. We find a similar albeit weaker signal between the progressive and conservative communities, confirming the role of left-leaning discourse in shaping the general terms of debate (?). We also find that Black celebrities significantly shape the discourse, introducing word changes that the center/left news media as well as the conservative community later take up. The scale of our data, combined with our methods of validation, provide a new layer of evidence to support the largely qualitative scholarship that affirms the role of the BLM movement for bringing ideas about racial justice from the margin to the mainstream. Our results also provide early evidence of the conservative backlash that is now being expressed in a range of “anti-woke” legislation and related policies and political groups, since we find that the conservative community is the largest follower of all word changes — not only after the summer of 2020 but from the movement’s very beginning.
@@
The topic of Twitter, and Black Twitter in particular, has been the subject of significant research in the fields of communication and media studies, among other humanities and social science fields. In Distributed Blackness (?), André Brock theorizes the community structure of Black Twitter as consisting of the people who are online and in conversation with each other at any given time. He also theorizes “Twitter time” as non-linear, defined more by specific events and the bursts of conversation they prompt than by any standard timescale. In Hashtag Activism (?), Sarah J. Jackson, Moya Bailey, and Brooke Foucault Welles propose a similar conception of Twitter communities organized by the hashtag itself. Hashtags, they explain, ”designate collective thoughts, ideas, arguments, and experiences that might otherwise stand alone.” Additional work on Black Twitter has considered the perspectives of the users themselves (?), its historical antecedents (?), and its possible futures (?). This qualitative work constitutes the foundation for our project. Our contribution represents an attempt to formalize the concepts and structures these scholars describe.
@@
Several characteristics of online social movements, including the motivation and role of participants (???), activist influence on potential recruits (?), formation and sustenance of collective identities (?),and mobilization of activists (?????), have been extensively studied. BLM, in particular, has been studied as a unique example of a highly contested social movement with a sizeable online footprint (????). The online discourse of this movement has been shaped by tragic offline events (?) and has persisted for almost a decade (?). In terms of our project, the closest related work is Freelon, McIlwain, and Clark’s ”Beyond the Hashtags” report (?), which employs community detection followed by hand-labeling by domain experts in order to trace the changes in network structure over the first 18 months of the BLM movement. Recognizing the expertise involved in that project, our goal was to carry forward the authors’ work. Here, we extend the timeframe of analysis by another five years; develop new methods that allow us to expand the initial set of communities labeled by that team; and introduce an additional layer of linguistic analysis that complements the original study.
Linguistic methods have also been used to study different aspects of BLM. For example, the dynamics of participation have been shown to be correlated with textual features of tweets (?); the impact of police killings has been shown to reflect in the emotional narratives of social media participants (?); hashtag use has been shown to draw attention (??) or push back against issues (?). Framing and rhetorical strategies have been shown to be means of differentiation and coalition forming (???). We extend this line of work by inducing the leadership structure within the coalition of communities through the adoption of linguistic changes by community members.
Table 1: Descriptive statistics User and tweet statistics for each temporal partition are provided. Start and end dates are inclusive for each period.
@@ 7 Aug 07, 2020 - Dec 31, 2020 146 5,112,396 877,086 214,050
Sociocultural changes often get encoded as lexical semantic changes in language (?). Consequently, computational methods to automatically extract semantic changes, given a timestamped text collection, have been developed with increasing regularity (??????, interalia). Semantic changes detected by such methods are demonstrably effective in two ways: 1) as first-order objects of analytical interest, such as in tracking the evolution of political issues (?) or shift in racial attitudes over time (?); 2) as units that can be aggregated to recover the latent leadership structure among scholarly articles (??) or newspapers (?). In this work, we follow the latter approach, using word embeddings to unveil the leadership dynamics between communities and their role in shaping online discourse of the BLM movement.
Figure 2: Volume of tweets per month We visualize the temporal distribution of tweets in our dataset in the style of Freelon et al (?). The seven temporal partitions are indicated here both by color and a reference label (e.g., Period 1). Some pivotal events to the BLM movement in each partition are annotated.
@@
Our dataset consists of tweets containing ”#BlackLivesMatter” between June 1, 2014 and December 31, 2020. These tweets were collected in two phases: tweets between June 1, 2014 and May 31, 2015 were rehydrated in April 2021 using twarc [^1] https://twarc-project.readthedocs.io/en/latest/ from tweet IDs provided by Freelon et al. ( ?). [^2] In addition to the #BlackLivesMatter hashtag, these tweets also included 44 additional ”keywords related to BLM and police killings of Black people under questionable circumstances” (?, p. 21) Our choice of this data source was motivated by our desire to build upon the hand-labeled communities in that paper, which we then used to seed our own community detection step (see next section). Tweets between June 1, 2015 and December 31, 2020, were collected using Twitter’s search API between April and August 2021. Together, this resulted in a total of 82,661,364 tweets from 14,914,359 unique users.
@@
Our analysis proceeds in two steps: a community detection step (see Community Detection and Matching) and a semantic leadership network induction step (see Leadership Network Induction). For the former, we filtered out all tweets tagged as non-English using langid (?); for the latter, we did additional filtering passes to remove tweets that were retweets, exact duplicates, or tagged as non-English by langid (?) and fasttext (?). The first round of filtering left 69,560,572 tweets; the second round of filtering left 32,652,123 tweets. The breakdown of tweets by timebin and unique users is shown in Table  1.
@@
We remain attentive to concerns about extracting knowledge and intellectual labor from online communities for academic research (?????) throughout our analysis. These concerns motivate our modeling approach, which is designed to extend rather than replace existing scholarship. In addition, our multiracial project team has members who have engaged with the BLM movement in various roles, on and offline. With respect to concerns about privacy and attribution, we are not required to identify individual users in most cases, as the majority of our analysis is undertaken at an aggregate level. In the four instances where we identify individual users or their tweets, we take a context-specific approach (?). 6 requires that we give a sense of the users that comprise each community. To avoid unwanted exposure, we rank users by in-degree and name only those within the top five of each community that a) still have public accounts on X; and b) meet the ”reasonably public” threshold of 3000 or more followers as formulated by Freelon et al. (?). Table 2 requires that we provide examples of the word changes that our model detects. Here, we followed current best practice by paraphrasing tweets rather than quoting directly (??) because we do not want to expose activist users or users with low follower counts to unwanted attention or possible harm (?). We also confirm via X that our ”ethically fabricated” language (?) is not traceable back to the original user. The two users mentioned by name in the paper are public figures and as such, have upwards of one million followers.
@@
We partition the dataset around the local maxima of tweet frequency, allowing irregular intervals. We find that the relative maxima correspond to real-world events relating to BLM. For instance, the first local maximum is found on August 11th, 2014, two days after the murder of Michael Brown  (??). We used these real-world events as a qualitative baseline for validating possible maxima. When additional maxima occur within 30 days of another maximum, they are grouped into one period (see Table 3). From the 11 relative maxima that were extracted using this method, six became markers for time periods resulting in seven intervals (see Figure 2 and Table 1).
@@
To obtain user clusters, we maximize intra-group modularity (??) by modifying the Louvain clustering algorithm (?) — a well-known algorithm for detection of an unconstrained number of communities — using the TSM package. Because of our goal of extending the communities identified in Freelon, McIlwain, and Clark (?), which were hand-labeled by experts on BLM and Black Twitter, we augment the basic Louvain approach with six seed communities, consisting of 60 users, which were reconstructed from the communities documented in the original study. This improved the overall coherence of the clusters, which initial experiments had demonstrated to be poor.
@@
Figure 3: Temporal tweet distribution The timeseries of tweet volume per community after filtering.
Figure 4: Community size Count of unique users in each community. Bar segments represent subcommunity clusters with height proportional to cluster size.
@@
To find persistent communities from the ephemeral clusters obtained in the previous step, we cast the problem as a matching problem in a bipartite graph. Specifically, for a pair of successive intervals, we construct a bipartite graph linking each cluster from one interval to every cluster in the next interval. The weight on these links represents similarity between the two clusters, and is calculated as a weighted Jaccard score taking into account the in-degree of the users in a cluster (?). Following Freelon et al., we experimented with pruning the “weaker” edges if their similarity was below 0.3. However, since our intervals span over months or years, we see a greater level of fragmentation at this threshold. Consequently, we used a knee plot, which highlights the tradeoff between the number of matched communities and the jaccard threshold. Using this diagnostic, we lower the threshold to 0.07, a brightline that permits persistent communities with noise confined to spurious matches (see Figure 7).
@@
To develop a stronger sense of each community as a whole, the annotators used the set of labels assigned to its top 20 users to define a short description of each community. Using the two to four alternative descriptions for each community, the annotators then agreed on a final short description. In order to conduct semantic change analysis and perform permutation tests, it is necessary to condense the number of communities analyzed. Using the final short descriptions, we merged our 65 communities into 12 larger groups (see Figure  4 and Figure  6; ”miscellaneous” not shown).
We excluded two groups from this analysis, one for not being in English and the other for containing spam. Since communities on Twitter change in membership over time, it is important to check if these group labels are apt. Once grouped into the 12 macro communities, the average in-degree of the top 20 users was no less than 20%, suggesting that, relative to the size of their respective communities, these users play a central role in its dialogue. Thus, their presence provides important information about the background of other users and potential insight into how they may approach discussing social issues.
@@
We adopted the method from ? (?) to induce a semantic leadership network between document sources. Soni et al. considered a set of historical newspapers as document sources and individual articles as documents. In contrast, we consider our Twitter user communities as document sources and individual tweets as documents.
Besides this superficial difference in the setup, there are two substantive differences. First, Soni et al. divided their timestamped collection of newspaper articles into temporal slices of equal timespans. In contrast, our tweet collection is divided into intervals demarcated by timestamps of real-world incidents. Second, Soni et al. considered document sources to remain fixed over time. In contrast, we consider communities of Twitter users as document sources which may undergo changes as a result of users joining, leaving and migrating between communities. We thus apply the methodology to our data but note that the interpretation of influence across communities should be qualified to account for these user dynamics.
@@
To identify semantic changes, we learn temporal word-type embeddings from data. Other methods to detect semantic changes using token embeddings have been proposed but with mixed success (?). Consequently, we approach the task as learning word embeddings and then enhancing them to account for information from temporal or other facets. Words are mapped to embeddings, typically by maximizing the following probability under the skipgram language model (?) between any pair of nearby tokens $(w_{j},w_{j^{\prime}})$ in a single document ,
@@
where $\mathbf{b}_{w_{i,j}}$ is the base embedding of the word $w_{i,j}$ and $\mathbf{r}_{w_{i,j}}^{(t_{i})}$ is the residual for time $t_{i}$. The residual turns the core meaning to a temporally specific meaning and are hence regularized towards zero to follow the assumption that for most words the core meaning remains intact over time. Embedding decomposition of this type has been used in other applications such as to identify geographic variation (?) or perceptual differences in meaning (?).
Once the base and the residual components of the input embeddings for each word are obtained, semantic changes can be found by comparing near-neighbors in the embedding space (?). The upshot, at the end of this step, is a ranked order list of triples, consisting of the word that has changed, the timestamp for the onset of the change, and the timestamp for the conclusion of the change.

### v2

@@ Abstract
In the wake of the 2024 US presidential election, pundits on both the left and the right pointed to a conservative backlash against “woke politics” to explain the election’s outcome. These politics, rooted in substantive beliefs about equity and justice–and particularly racial justice–owe their most recent rise to prominence to the Black Lives Matter (BLM) movement. A significant body of work, both qualitative and quantitative, has documented how BLM was able to move these beliefs from the margin to the mainstream. In this paper, we focus on the words that index these beliefs, devising a novel method of modeling semantic leadership across a set of communities associated with the BLM movement that is informed by domain-specific theory about Black Twitter. We describe our bespoke approaches to time-binning, community clustering, and connecting communities over time, as well as our adaptation of state-of-the-art approaches to semantic change detection and semantic leadership induction. We find evidence at scale of the leadership role of BLM activists and progressives, as well as of Black celebrities. We also find evidence of sustained conservative engagement with this discourse, suggesting an alternative explanation for how we have arrived at the present political moment.
@@ Introduction
In early 2020, when the Black Lives Matter (BLM) movement had secured its place in the US national consciousness and when this project began, it was still necessary to provide evidence of how battles over social and political change are waged through both words and action. Writing again in April 2025, as we prepare this paper for final submission, it is evident to all (or should be) that words index larger concepts and debates. We need look no further than the US federal government, where ideas about equity and justice–and racial justice in particular–have not only become renewed topics of debate; the terms “equity” and “racial justice” themselves have been banned from government use (Yourish et al. 2025). While this act of censorship serves as a negative example, ample evidence also points to how words can open up conversations, introduce new ideas into public consciousness, and positively impact social behavior and government policy alike (Dunivin et al. 2022; Yan, Chiang, and Lin 2024). In this paper, we return to the Black Lives Matter movement, which we see as one of the most successful contemporary examples of the positive value of words (among other contributions), to explore with precision how words enter into a specific sociopolitical conversation, how they transform as that conversation evolves, and who is responsible for carrying forward those new or changed meanings.
@@
Figure 1: Semantic leadership in #BlackLivesMatter: Nodes represent the most central communities of the BLM network. The thickness of the edges are proportional to the number of word changes shared by each pair.
@@
While there exists abundant research on Black Lives Matter (BLM), both as a hashtag (Choudhury et al. 2016; Jackson, Bailey, and Welles 2020; Le-Khac, Antoniak, and So 2023; Giorgi et al. 2022; Jones, Nurse, and Li 2022) and a movement (Freelon, McIlwain, and Clark 2016; del Rio 2020; Clark 2024), there are still open questions about how precisely BLM was able to “push the mainstream public sphere on issues of social progress” (Jackson, Bailey, and Welles 2020, p. xxvii). This is largely due to the unique community structure of Twitter, in which communities are defined not by individual users but by hashtags, and by the conversations among users that hashtags enable (Tufekci 2013; Spiro and Monroy-Hernández 2016; Jackson, Bailey, and Welles 2020; Brock 2020). Operationalizing these medium-specific theories about community formation on Twitter, as well as about “Twitter time,” we present a multipart model that allows us to 1) detect the latent communities associated with the #BlackLivesMatter hashtag and link them over time; and 2) detect semantic changes associated with each community using word embeddings. Together with a measure of semantic leadership (Soni, Klein, and Eisenstein 2021), we are able to induce a semantic leadership network (see Figure 1), allowing us to identify the communities responsible for introducing new or changed word meanings into the BLM network and the communities responsible for adopting (or co-opting) them.
@@
We find that the community of BLM activists plays an outsized role in introducing semantic changes into the network, and that these changes are most consistently taken up by the conservative community. We find a similar albeit weaker signal between the progressive and conservative communities, confirming the role of left-leaning discourse in shaping the general terms of debate (Le-Khac, Antoniak, and So 2023). We also find that Black celebrities significantly shape the discourse, introducing word changes that the center/left news media as well as the conservative community later take up. The scale of our data, combined with our methods of validation, provide a new layer of evidence to support the largely qualitative scholarship that affirms the role of the BLM movement for bringing ideas about racial justice from the margin to the mainstream. Our results also provide an alternative explanation for how we have arrived at the present political moment. Contrary to a narrative of conservative backlash against a movement that became “too woke”, we find that the conservative community — as the largest follower of word changes across the entire timespan of our study — had been engaging directly with the BLM movement from its very start. Today, we see the end goal of this engagement: to distort and disrupt the messaging of the BLM movement, and ultimately, to weaponize the movement’s words against its most closely held beliefs.
@@
Figure 2: Volume of tweets per month: We visualize the temporal distribution of tweets in our dataset in the style of Freelon et al (2016). The seven temporal partitions are indicated here both by color and a reference label (e.g., Period 1). Some pivotal events to the BLM movement in each partition are annotated.
@@
The topic of Twitter, and Black Twitter in particular, has been the subject of significant research in the fields of communication and media studies, among other humanities and social science fields. In Distributed Blackness (2020), André Brock theorizes the community structure of Black Twitter as consisting of the people who are online and in conversation with each other at any given time. He also theorizes “Twitter time” as non-linear, defined more by specific events and the bursts of conversation they prompt than by any standard timescale. In Hashtag Activism (2020), Sarah J. Jackson, Moya Bailey, and Brooke Foucault Welles propose a similar conception of Twitter communities organized by the hashtag itself. Hashtags, they explain, “designate collective thoughts, ideas, arguments, and experiences that might otherwise stand alone.” Additional work on Black Twitter has considered the perspectives of the users themselves (Clark 2024), its historical antecedents (Klassen et al. 2022), and its possible futures (Walcott 2024). This qualitative work constitutes the foundation for our project. Our contribution represents an attempt to formalize the concepts and structures that these media studies scholars describe.
@@
Several characteristics of online social movements, including the motivation and role of participants (Tufekci and Wilson 2012; Bozarth and Budak 2017; Jones, Nurse, and Li 2022), activist influence on potential recruits (González-Bailón et al. 2011), formation and sustenance of collective identities (Brown et al. 2017), and mobilization of activists (Theocharis et al. 2015; Freelon, McIlwain, and Clark 2018; Mundt, Ross, and Burnett 2018; Yan, Chiang, and Lin 2024; Mendelsohn et al. 2024), have been extensively studied. BLM, in particular, has been studied as a unique example of a highly contested social movement with a sizeable online footprint (Arif, Stewart, and Starbird 2018; Stewart et al. 2017; Gallagher et al. 2018; Giorgi et al. 2022). The online discourse of this movement has been shaped by tragic offline events (Peng, Budak, and Romero 2019) and has persisted for almost a decade (Dunivin et al. 2022). In terms of our project, the closest related work is Freelon, McIlwain, and Clark’s “Beyond the Hashtags” report (2016), which employs community detection followed by hand-labeling by domain experts in order to trace the changes in network structure over the first 18 months of the BLM movement. Recognizing the knowledge and expertise involved in this project, our goal was to carry forward the authors’ work. Here, we extend the timeframe of analysis by another five years; develop new methods that allow us to expand the initial set of communities labeled by their team; and introduce an additional layer of linguistic analysis that complements the original study.
Linguistic methods have also been used to study different aspects of BLM. For example, the dynamics of participation have been shown to be correlated with textual features of tweets (Choudhury et al. 2016); the impact of police killings has been shown to reflect in the emotional narratives of social media participants (Field et al. 2022); hashtag use has been shown to draw attention (Blevins et al. 2019; Yan, Chiang, and Lin 2024) or push back against issues (Gallagher et al. 2018). Framing and rhetorical strategies have been shown to be means of differentiation and coalition forming (Stewart et al. 2017; Wilkins, Livingstone, and Levine 2019; Mendelsohn et al. 2024). We extend this line of work by inducing the leadership structure within the coalition of communities through the adoption of linguistic changes by community members.
@@ 7 Aug 07, 2020 - Dec 31, 2020 146 5,112,396 877,086 214,050
Table 1: Descriptive statistics: Tweet and user statistics for each temporal partition. Start and end dates are inclusive.
@@
Sociocultural changes often get encoded as lexical semantic changes in language (Tahmasebi, Borin, and Jatowt 2019). Consequently, computational methods to automatically extract semantic changes, given a timestamped text collection, have been developed with increasing regularity (Wijaya and Yeniterzi 2011; Kim et al. 2014; Kulkarni et al. 2015; Hamilton, Leskovec, and Jurafsky 2016b; Giulianelli, Del Tredici, and Fernández 2020; Card 2023, interalia). Semantic changes detected by such methods are demonstrably effective in two ways: 1) as first-order objects of analytical interest, such as in tracking the evolution of political issues (Rudolph and Blei 2018) or shift in racial attitudes over time (Garg et al. 2018); 2) as units that can be aggregated to recover the latent leadership structure among scholarly articles (Soni, Lerman, and Eisenstein 2020; Soni, Bamman, and Eisenstein 2022) or newspapers (Soni, Klein, and Eisenstein 2021). In this work, we follow the latter approach, using word embeddings to unveil the leadership dynamics between communities and their role in shaping online discourse of the BLM movement.
@@
Our dataset consists of tweets containing “#BlackLivesMatter” between June 1, 2014 and December 31, 2020. These tweets were collected in two phases: tweets between June 1, 2014 and May 31, 2015 were rehydrated in April 2021 using twarc [^1] https://twarc-project.readthedocs.io/en/latest/ from tweet IDs provided by Freelon et al. ( 2016). [^2] In addition to the #BlackLivesMatter hashtag, these tweets also included 44 additional “keywords related to BLM and police killings of Black people under questionable circumstances” (Freelon, McIlwain, and Clark 2016, p. 21) Our choice of this data source was motivated by our desire to build upon the hand-labeled communities in that paper, which we then used to seed our own community detection step (see next section). Tweets between June 1, 2015 and December 31, 2020, were collected using Twitter’s search API between April and August 2021. Together, this resulted in a total of 82,661,364 tweets from 14,914,359 unique users.
@@
Our analysis proceeds in two steps: a community detection step (see Community Detection and Matching) and a semantic leadership network induction step (see Leadership Network Induction). For the former, we filtered out all tweets tagged as non-English using langid (Lui and Baldwin 2012); for the latter, we did additional filtering passes to remove tweets that were retweets, exact duplicates, or tagged as non-English by langid (Lui and Baldwin 2012) and fasttext (Grave et al. 2018). The first round of filtering left 69,560,572 tweets; the second round of filtering left 32,652,123 tweets. The breakdown of tweets by timebin and unique users is shown in Table 1.
@@
Throughout our analysis, we remain attentive to concerns about extracting knowledge and intellectual labor from online communities for academic research (Bruckman 2002; Jules, Summers, and Mitchell 2018; Dym and Fiesler 2020; Jackson, Bailey, and Welles 2020; Walsh 2023). These concerns motivate our modeling approach, which is designed to extend rather than replace existing scholarship. In addition, our multiracial project team has members who have engaged with the BLM movement in various roles, on and offline. With respect to concerns about privacy and attribution, we are not required to identify individual users in most cases, as the majority of our analysis is undertaken at an aggregate level. In the four instances where we identify individual users or their tweets, we take a context-specific approach (Dym and Fiesler 2020). Figure 6 requires that we give a sense of the users that comprise each community. In this case, to avoid unwanted exposure, we rank users by in-degree and name only those within the top five of each community that a) have maintained public accounts through Twitter’s transition to X; and b) as of September 2024, still meet the “reasonably public” threshold of 3000 or more followers as formulated by Freelon et al. (2016). Table 2 requires that we provide examples of the word changes that our model detects. Here, we follow current best practice by paraphrasing tweets rather than quoting directly (Dym and Fiesler 2020; Walsh 2024) because we do not want to expose activist users or users with low follower counts to unwanted attention or possible harm (Jules, Summers, and Mitchell 2018). We also confirm via the current X.com search interface that our “ethically fabricated” language (Markham 2012) is not traceable back to the original user. The two users mentioned by name in the paper are public figures and as such, have upwards of one million followers.
@@
We partition the dataset around the local maxima of tweet frequency, allowing irregular intervals. We find that the relative maxima correspond to real-world events relating to BLM. For instance, the first local maximum is found on August 11th, 2014, two days after the murder of Michael Brown  (Anderson 2016; Freelon, McIlwain, and Clark 2016). We use these real-world events as a qualitative baseline for validating possible maxima. When additional maxima occur within 30 days of another maximum, they are grouped into one period (see Table 3). From the 11 relative maxima that were extracted using this method, six became markers for time periods resulting in seven intervals (see Figure 2 and Table 1).
@@
To obtain user clusters, we maximize intra-group modularity (Newman and Girvan 2004; Newman 2006) by modifying the Louvain clustering algorithm (Blondel et al. 2008) — a well-known algorithm for detection of an unconstrained number of communities — using the TSM package.
Because of our goal of extending the communities identified in Freelon, McIlwain, and Clark (2016), which were hand-labeled by experts on BLM and Black Twitter, we augment the basic Louvain approach with six seed communities, consisting of 60 users, which were reconstructed from the communities documented in the original study. This improved the overall coherence of the clusters, which initial experiments had demonstrated to be poor.
@@
Figure 3: Tweet counts over time after filtering, by community. Grouping subcommunities allows for more consistent representation of each community over the entire span.
Figure 4: Community size: Count of unique users in each community. Bar segments represent subcommunity clusters with height proportional to cluster size.
@@
To find persistent communities from the ephemeral clusters obtained in the previous step, we cast the problem as a matching problem in a bipartite graph. Specifically, for a pair of successive intervals, we construct a bipartite graph linking each cluster from one interval to every cluster in the next interval. The weight on these links represents similarity between the two clusters, and is calculated as a weighted Jaccard score taking into account the in-degree of the users in a cluster (Freelon, Lynch, and Aday 2015). Following Freelon et al., we experimented with pruning the “weaker” edges if their similarity was below 0.3. However, since our intervals span over months or years, we see a greater level of fragmentation at this threshold. Consequently, we used a knee plot, which highlights the tradeoff between the number of matched communities and the jaccard threshold. Using this diagnostic, we lower the threshold to 0.07, a brightline that permits persistent communities with noise confined to spurious matches (see Figure 7).
@@
To develop a stronger sense of each community as a whole, the annotators used the set of labels assigned to its top 20 users to define a short description of each community. Using the two to four alternative descriptions for each community, the annotators then agreed on a final short description. In order to conduct semantic change analysis and perform permutation tests, it is necessary to condense the number of communities analyzed. Using the final short descriptions, we merged our 65 communities into 12 larger groups (see Figure  4 and Figure  6; “miscellaneous” not shown).
We excluded two groups from this analysis, one for not being written in English and the other for containing spam. Since communities on Twitter change in membership over time, it is important to check if these group labels are apt. Once grouped into the 12 macro communities, the average in-degree of the top 20 users was no less than 20%, suggesting that, relative to the size of their respective communities, these users play a central role in its dialogue. Thus, their presence provides important information about the background of other users and potential insight into how they may approach discussing social issues.
@@
We adopted the method from Soni, Klein, and Eisenstein (2021) to induce a semantic leadership network between document sources. Soni et al. 2021 considered a set of historical newspapers as document sources and individual articles as documents. In contrast, we consider our Twitter user communities as document sources and individual tweets as documents.
Besides this superficial difference in the setup, there are two substantive differences. First, Soni et al. 2021 divided their timestamped collection of newspaper articles into temporal slices of equal timespans. In contrast, our tweet collection is divided into intervals demarcated by timestamps of real-world incidents. Second, Soni et al. 2021 considered document sources to remain fixed over time. In contrast, we consider communities of Twitter users as document sources which may undergo changes as a result of users joining, leaving and migrating between communities. We thus apply the methodology to our data but note that the interpretation of influence across communities should be qualified to account for these user dynamics.
@@
To identify semantic changes, we learn temporal word-type embeddings from data. Other methods to detect semantic changes using token embeddings have been proposed but with mixed success (Laicher et al. 2020). Consequently, we approach the task as learning word embeddings and then enhancing them to account for information from temporal or other facets. Words are mapped to embeddings, typically by maximizing the following probability under the skipgram language model (Mikolov et al. 2013) between any pair of nearby tokens $(w_{j},w_{j^{\prime}})$ in a single document ,
@@
where $\mathbf{b}_{w_{i,j}}$ is the base embedding of the word $w_{i,j}$ and $\mathbf{r}_{w_{i,j}}^{(t_{i})}$ is the residual for time $t_{i}$. The residual turns the core meaning to a temporally specific meaning and are hence regularized towards zero to follow the assumption that for most words the core meaning remains intact over time. Embedding decomposition of this type has been used in other applications such as to identify geographic variation (Bamman, Dyer, and Smith 2014) or perceptual differences in meaning (Gillani and Levy 2019).
Once the base and the residual components of the input embeddings for each word are obtained, semantic changes can be found by comparing near-neighbors in the embedding space (Hamilton, Leskovec, and Jurafsky 2016a). The upshot, at the end of this step, is a ranked order list of triples, consisting of the word that has changed, the timestamp for the onset of the change, and the timestamp for the conclusion of the change.